STRUCTURAL STUDIES OF PYRROLE-BENZENE COMPLEXES BY CHIRPED-PULSE ROTATIONAL SPECTROSCOPY

\underline{SIMON LOBSIGER}, CRISTOBAL PEREZ, DANIEL P. ZALESKI, NATHAN A. SEIFERT, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904; CHANTAL PFAFFEN, MARIA A. TRACHSEL, SAMUEL LEUTWYLER, Departement für Chemie und Biochemie, Universität Bern, Freiestrasse 3, 3012 Bern, Switzerland.

Non-covalent intermolecular interactions are important in structural biology. The N-H ⋯ π hydrogen bond between amino acid side chains is an important structural determinant and highly affects the secondary structure of proteins. The pyrrole-benzene complex can be viewed as a model system for studying these fundamental interactions. Previous IR and UV spectroscopic studies of the pyrrole-benzene complex by Dauster et al.^{a} and Pfaffen et al.^{b} support a T-shaped structure with an N-H ⋯ π hydrogen bond to the benzene ring. In order to obtain accurate structural information we have investigated the broadband rotational spectrum of the supersonic-jet cooled complexes of pyrrole with benzene and benzene-d_{1} in the 2-18 GHz frequency range. In addition to the hetero dimer we have also observed the two cyclic mixed trimers (pyrrole)_{2}-benzene and pyrrole-(benzene)_{2}.

aI. Dauster, C. A. Rice, P. Zielke, and M. A. Suhm Phys. Chem. Chem. Phys. \underline{10}, 2827 (2008)

bC. Pfaffen, D. Infanger, P. Ottiger, H. M. Frey, and S. Leutwyler Phys. Chem. Chem. Phys. \underline{13}, 14110 (2011)